Probability of a customer making a purchase is 0.7. Four customers walk into the store. What is the probability that two of the four customers will not make a purchase?


1
Expert's answer
2021-08-23T09:07:03-0400

"P(X=x) = C^n_xp^x(1-p)^{n-x} \\

p=0.7 \\

n= 4"

Let X= number of customers making a purchase.

"P(X=2) = C^4_2 \times 0.7^2 \times (1-0.7)^{4-2} \\

= \frac{4!}{2!(4-2)!} \times 0.49 \times 0.09 \\

= 6 \times 0.49 \times 0.09 \\

= 0.2646 \\

= 26.46 \; \%"

The probability that two of the four customers will not make a purchase is 26.46 %.